1/10 1/5 3/10 2/5 1/2 3/5 a,1/4 b,4/5 c,7/10 d,8/10
answer of this number series question

Found 3 solutions by CubeyThePenguin, ikleyn, greenestamps:
Answer by CubeyThePenguin(3113)

1/10, 1/5, 3/10, 2/5, 1/2, 3/5, ___

next in the sequence = 3/5 + 1/10 = 7/10

Answer by ikleyn(53762)

.

Notice that this post came with no question.


Such posts are PROHIBITED at this forum.


They are considered as nonsensical and are the subject of ignoring.



Answer by greenestamps(13334)



Write all the numbers with a common denominator:

1/10, 2/10, 3/10, 4/10, 5/10, 6/10, _____

Clearly the intended answer is 7/10.